### v1

# Unleashing the Power of Large Language Models for Group POI Recommendations

Abstract.
Group Point-of-Interest (POI) recommendations aim to predict the next POI that satisfies the diverse preferences of a group of users. This task is more challenging than traditional individual POI recommendations due to complex group decision-making and extremely sparse group-level check-in data. Existing methods for group POI recommendations primarily rely on single ID-based features from check-in data, capturing only statistical correlations and failing to fully utilize the rich semantic information contained in the check-ins, resulting in suboptimal performance. To this end, we propose a framework that unleashes the power of the Large Language Model (LLM) for context-aware group POI recommendations (LLMGPR). Our approach first introduces POI tokens alongside the original word tokens of the LLM, which are initialized by applying the LLM to the rich information of each POI. We then propose a novel sequencing adapter guided by Quantized Low-Rank Adaptation (QLORA) to modify the LLM. The enhanced LLM can learn sequence representations by combining semantic-enhanced POI tokens and rich contextual information including positional encodings and spatio-temporal differences. This approach can be adapted for learning either group or user representations depending on the sequence type. Furthermore, we enhance group representations by aggregating individual member representations with another QLORA-based aggregation adapter and introducing a self-supervised learning task that predicts the purpose of check-in sequences, alleviating the data sparsity issue. Our experimental results demonstrate that LLMGPR outperforms existing methods, effectively addressing group-level data sparsity and providing superior recommendations.
*Corresponding author
copyright: none
1. Introduction
With the rapid expansion of Location-Based Social Networks (LBSNs) such as Foursquare and Weeplace, Point-of-Interest (POI) recommendations have become essential for helping individuals discover interesting places. These recommendations are crucial for applications like mobility prediction, route planning, and location-based advertising (Li et al., 2018). While individual POI recommendation systems are effective in addressing personal preferences by leveraging users’ check-in histories, they are limited in scenarios involving group decision-making. As location-based social activities grow, there is a distinct need for recommendations that consider the preferences of multiple users simultaneously, known as Group POI Recommendation. Group-level recommendations are essential in various scenarios, such as friends planning a weekend outing, families on vacation, or colleagues on a business trip, where meeting the collective preferences of all group members is crucial.
Figure 1. (a) Example of individual POI recommendation. (b) Example of group POI recommendation.
Group POI recommendation systems typically involve a two-phase pipeline. In the first phase, the group is treated as a virtual user, allowing traditional individual POI recommendation methods, such as matrix factorization (Lian et al., 2014), transformer-based approaches (Luo et al., 2021), graph-based approaches (Wang et al., 2022), and diffusion-based approaches (Qin et al., 2023), to be directly applied. These methods learn group representations based on group-level check-in data and individual representations based on individual-level check-in data, respectively. In the second phase, these group representations are further enhanced by aggregating the individuals’ representations within the group. Aggregation approaches include the simple average (Vinh Tran et al., 2019), weighted average (Sojahrood and Taleai, 2021), attention mechanism (Sankar et al., 2020), and graph convolution network (Li and Zhu, 2024). Despite their effectiveness, current group POI recommendations typically depend only on the POI-ID sequence data ( including individual and group check-in sequences). This means the models can only capture statistical information from check-ins, while the rich semantic information inherent in the data remains underutilized.
Benefiting from the recent breakthroughs in pre-trained large language models (LLMs), which have significantly advanced the semantic understanding of natural language, some research has started to explore integrating LLMs into POI recommendation systems to better capture the rich semantic information contained in POI data (e.g., POI categories.), thereby alleviating data sparsity and cold-start issues. For example, LLMMove (Feng et al., 2024) and LLM-Mob (Wang et al., 2023) utilize a series of carefully designed prompts to input individual users’ check-in records into pre-trained LLMs, leveraging the commonsense knowledge embedded in the LLMs to better exploit the rich semantic information in check-in data. Additionally, LLM4POI (Li et al., 2024) further fine-tunes the pre-trained LLM using individual check-in data to enhance its performance.
However, current LLM-based POI recommendation methods are primarily designed for individual POI recommendations, making them non-trivial to work for the more challenging group POI recommendations due to the following challenges: (1) Diverse Preferences and Complex Group Decision-making: Unlike individual POI recommendations that only need to satisfy the preferences of a single user, as shown in Figure 1, group POI recommendations must find a consensus among group members on a POI to visit. However, the interests and preferences of individuals within a group are often diverse, and sometimes a group’s preferences are even different from the preferences of each group member (Guo et al., 2021), making it challenging to model the complex process of group decision-making. Existing group POI recommendation works (Liu et al., 2024; Li et al., 2020) are mainly based on the POI-ID sequence data, failing to capture the complex factors influencing group decision-making. On the other hand, existing LLM-based individual POI recommendation methods are typically designed to handle the preferences of a single user, leveraging personal check-in records and profile data. These methods do not account for the diverse individual preferences in a group and the complex group decision-making, which is essential for effective group recommendations. (2) Extreme Data Sparsity: Unlike individual check-ins, which are recorded by single users visiting a POI, group check-in data involves multiple users checking in to the same location together, such as a group of friends checking in together at a restaurant. Group check-ins happen less frequently because they require coordination among several users, making group check-in data significantly sparser than individual check-in data. For example, as discussed in Section 4.1, the group check-in data in two public datasets is approximately 50-100 times sparser than individual check-in data. Therefore, despite the fact that existing LLM-based POI recommendation methods have shown potential in mitigating individual data sparsity by utilizing the semantic information in check-in data, they face greater challenges when dealing with group check-in data. The extremely limited amount of group check-ins makes it difficult to fine-tune models using the sparse group check-in records.
In this work, we introduce a pre-trained large language model-based framework for group POI recommendations (LLMGPR). Specifically, to address the challenge of Diverse Preferences and Complex Group Decision-making, our framework first introduces POI tokens alongside the original word tokens of the LLM. POI embeddings are initialized by applying the LLM to the rich information of each POI (e.g., category and address). By combining a semantic-enhanced POI token and rich contextual information (i.e., positional encodings and spatio-temporal differences) to model each check-in activity, the LLM can capture the complex and diverse factors for the group decision-making beyond the aggregation of individual preferences, such as sequential effect, geographical influence, temporal effect, and semantic effect (Xie et al., 2016). Then, based on the representation of each check-in activity, we propose a novel sequencing adapter under the guidance of Quantized Low-Rank Adaptation (QLORA) to learn the representation of a check-in sequence, which can be used as either group representation or user representation depending on the check-in sequence types.
To address the Extreme Data Sparsity challenge, we first propose a novel aggregation adapter to aggregate individual member representations obtained from their check-in sequences to enhance our group representations. Then, we introduce a novel self-supervised learning (SSL) task to further enhance group representations by learning and predicting sequence purposes (e.g., business trips and family vacations). Specifically, we first designed a prompt that leverages the spatio-temporal and contextual information within each sequence, enabling a pre-trained large language model to assign a category reflecting its underlying purpose, which then serves as the ground truth for self-supervised learning. In this way, the SSL signal helps the LLM capture and utilize additional semantic information of sparse group-level check-in data to refine the group representations. By jointly optimizing the self-supervised learning task (i.e., pretext task) and the adapters for the recommendation task, our framework effectively leverages the rich semantic information in check-in data to provide accurate and robust group-level POI recommendations. In conclusion, the contributions of this work are summarized as follows:
  • We propose LLMGPR, a pre-trained LLM-based framework for the next group POI recommendation. To the best of our knowledge, this is the first work to integrate pre-trained LLMs in the context of group POI recommendations.
  • We introduce semantic-enhanced POI tokens and integrate rich contextual information to represent each group check-in activity, accounting for the complex and diverse factors influencing group decisions. Based on the representation of each check-in activity, we propose a novel sequencing adapter based on QLORA to learn the group representation from the group’s check-in sequence.
  • To address data sparsity of group-level check-in data, we propose two methods to enhance the group representation: we first propose a novel aggregation adapter to combine individual member representations and then introduce a novel self-supervised learning (SSL) task that predicts sequence purposes, empowering the model to obtain more relevant semantic information.
  • We conduct extensive experiments to demonstrate the effectiveness of the proposed LLMGPR, showing significant improvements in accuracy for group POI recommendations.
2. PRELIMINARIES
This section lists key notations used throughout this paper, outlines our primary task, and introduces Large Language Models and Quantized Low-Rank Adaptation as the background knowledge.
2.1. Notations
We denote the sets of users, POIs, categories, and groups as $\mathcal{U}$, $\mathcal{P}$, $\mathcal{C}$, and $\mathcal{G}$ respectively. Each POI $p\in\mathcal{P}$ is associated with a category tag (e.g., entertainment or restaurant) $c_{p}\in\mathcal{C}$ and coordinates $(lon_{p},lat_{p})$. In addition, each group $g_{n}\in\mathcal{G}$ contains $K$ members, denoted as $\mathcal{K}_{n}=\{u_{1},u_{2},...,u_{K}\}$.
Definition 1: Check-in Sequence. A check-in activity indicates a user $u\in\mathcal{U}$ or a group $g\in\mathcal{G}$ has visited a POI $p\in\mathcal{P}$ at the timestamp $t$. By sorting all check-ins of the user or the group chronologically, a check-in sequence is obtained which contains $M$ consecutive POIs, denoted by $\mathcal{X}_{u}=\{p_{1},p_{2},...,p_{M}\}$ for the user $u$ or $\mathcal{X}_{g}=\{p_{1},p_{2},...,p_{M}\}$ for the group $g$.
2.2. Task: Group POI Recommendation
Given a group $g_{n}$, its check-in sequence $\mathcal{X}_{g_{n}}$, and individual check-in sequences of all group members $\mathcal{X}(\mathcal{K}_{n})=\{\mathcal{X}_{u_{1}},\mathcal{X}_{u_{2}},...,\mathcal{X}_{u_{K}}\}$, our goal is to generate a ranked list of possible POIs that group $g_{n}$ would like to visit next.
2.3. Large Language Models
Pre-trained large language models (LLMs) like GPT, T5, and Llama are advanced AI systems based on the Transformer architecture (Vaswani et al., 2017), which uses a self-attention mechanism to capture long-range dependencies in text. This work is based on Llama3-8b (Touvron et al., 2023), where input text sequences are tokenized, embedded, and processed through multiple Transformer blocks with layers for multi-head attention, feedforward transformations, and normalization. The model then generates tokens sequentially, producing coherent and contextually relevant text based on the given input and context.
2.4. Quantized Low-Rank Adaptation
Quantized Low-Rank Adaptation (QLoRA) (Dettmers et al., 2024) enhances LLM efficiency by quantizing model weights and applying low-rank adaptations for task-specific fine-tuning. Quantization reduces weight precision, lowering memory and computational needs. For a linear weight matrix W, the quantized version $\textbf{W}_{q}$ is:
(1) $\textbf{W}_{q}=round(\frac{\textbf{W}-min(\textbf{W})}{\Delta}),$
where $\Delta$ is the quantization step size:
(2) $\Delta=round(\frac{max(\textbf{W})-min(\textbf{W})}{2^{b}-1}),$
with $b$ being the quantization bits. Low-Rank Adaptation modifies $\textbf{W}_{q}\in\mathbb{R}^{d\times d}$ by:
(3) $\textbf{W}_{q}=\textbf{W}_{q}+\textbf{A}\cdot\textbf{B},$
where $\textbf{A}\in\mathbb{R}^{d\times r}$ and $\textbf{B}\in\mathbb{R}^{r\times d}$ are low-rank matrices updated during fine-tuning, and $r$ is the rank controlling adapter sizes. QLoRA is particularly useful for optimizing the LLM across tasks while conserving resources.
Figure 2. The overview of our proposed LLMGPR.
3. Methodology
In this section, we formally introduce the design of LLMGPR, with an overview provided in Figure 2. The main components include: (1) Learning group representation under the guidance of the LLM. (2) Enhancing group representation by leveraging the LLM to aggregate preferences of group members. (3) A self-supervised learning task to further enhance group representations.
3.1. Learning Group Representation
The proposed framework introduces a methodology for learning group representations based on group check-in sequences, which is crucial for capturing group preferences and making high-quality recommendations. Existing methods for modeling check-in sequences have the following limitations. Firstly, they often overlook the semantic information embedded in POI names and categories, failing to capture the sequential effect of check-in activities at the semantic level. Secondly, they ignore the spatial and temporal differences between check-in activities, missing out on capturing precise contextual information. For example, when there is a large spatial or temporal difference between two consecutive check-ins, the earlier check-in should not be considered as context for the later one. This is a key difference from word sequences. Thirdly, these methods usually require extensive retraining to adapt to new groups or regions, limiting their scalability and practical application.
To address these limitations, we introduce the large language model, which can leverage its world knowledge to enhance the POIs’ representations by exploiting their names and categories. By encoding temporal and spatial differences between check-in activities, the LLM captures the dynamic nature of user movements and preferences, resulting in richer and more context-aware representations. It excels at modeling sequential data, capturing long-range dependencies and contextual relationships between check-in activities. This allows for a more nuanced understanding of user behavior and improves prediction accuracy. Moreover, pre-trained on vast data from diverse domains, the LLM generalizes well to new users and regions without extensive retraining, making them more adaptable and scalable.
Before establishing the sequence representation semantically, we first introduce POI embeddings. While semantic representations capture the nuanced meanings of POI names and categories, relying solely on semantics may not fully leverage the unique identifiers of POIs. By incorporating POI embeddings through their IDs, we can ensure that each location is distinctly recognized and appropriately contextualized. This method allows us to directly reference specific POIs, enhancing the precision of recommendations and maintaining the integrity of location-specific data. Combining semantic information with POI embeddings provides a comprehensive approach that balances the depth of semantic understanding with the specificity of location identifiers. Formally, we use $\textbf{E}_{poi}\in\mathbb{R}^{P\times d}$ to denote the embedded representations of all POIs, where $P$ is the number of all POIs and $d$ is the embedding size which is the same as the size of the token embedding. Thus, all POIs can be regarded as additional tokens to the large language model. To maintain semantic information, instead of random initialization, the embedding of each POI is initialized by the large language model with a carefully designed prompt which is shown in Figure 3. Given the prompt, we directly use the output of the final transformer block as the initialized POI embedding without going through the linear and softmax layers. In addition, these embeddings will be further refined, making them adaptable for the task of group POI recommendations.
Figure 3. The prompt for initializing POI embeddings.
Given all POI embeddings, we design another prompt shown in Figure 4 to generate the representation of the check-in sequence. When this prompt is processed by the word embedding layer, the respective embedding in $\textbf{E}_{poi}$ will be utilized as the representation of the “POIID”, instead of treating it as a standard word. Similar to the initialization of POI embedding, the output of the final transformer block will be regarded as the embedded representation of the check-in sequence. This process is denoted as:
(4) $\textbf{e}_{g_{n}}=\Theta(\mathcal{X}_{g_{n}},\textbf{E}_{poi}),$
where $\textbf{e}_{g_{n}}\in\mathbb{R}^{d}$ is the embedding of the group check-in sequence $\mathcal{X}_{g_{n}}$, and $\Theta$ refers to the original large language model. Given the group embedding $\textbf{e}_{g_{n}}$ and the embeddings of $h$ candidate POIs $\textbf{E}_{cand}=[\textbf{e}_{p_{1}},\textbf{e}_{p_{2}},...,\textbf{e}_{p_{h}}]\in\mathbb{R}^{h\times d}$, the likelihood $\textbf{a}\in\mathbb{R}^{h}$ of candidate POIs to be visited next by this group is calculated as:
(5) $\textbf{a}=Softmax(\textbf{E}_{cand}\textbf{e}_{g_{n}}^{T}).$
Figure 4. The prompt for generating sequence embeddings.
Fine-tuning a large language model for POI recommendation, rather than relying on zero-shot learning, is essential because it tailors the model to the specific nuances and requirements of the task. This adaptation enhances relevance and accuracy by allowing the model to better understand domain-specific data and tasks. By integrating the specialized check-in sequence data, fine-tuning significantly improves the model’s performance, resulting in more accurate recommendations. Hence, we further employ the Quantized Low-Rank Adaptation (QLoRA) to fine-tune the LLM for the specific task of generating check-in sequence representations. Specifically, QLoRA is applied to all linear layers in $\Theta$, while the detailed approach is described in Section 2.4.
Here, we use $\Theta^{{}^{\prime}}_{s}$ to denote the adapter containing all additional weights from QLoRA for generating check-in sequence representations. During the fine-tuning process, the original LLM $\Theta$ is frozen while the sequencing adapter $\Theta^{{}^{\prime}}_{s}$ and the POI embedding matrix $\textbf{E}_{POI}$ are updated by the POI prediction objective:
(6) $\mathcal{L}_{poi}=\mathcal{L}_{CE}(y^{\prime},y),$
where ${L}_{CE}$ is the cross-entropy loss, $y^{\prime}$ is the predicted POI based on Equation 5, and $y$ is the ground truth POI. For a check-in sequence $\mathcal{X}$, the predictions are made successively on sequences $\{p_{1}\},\{p_{1},p_{2}\},...,\{p_{1},$ $p_{2},...,p_{M_{i}-1}\}$, and $\{p_{2},p_{3},...,p_{M_{i}}\}$ is the set of corresponding ground truth POIs. Due to the limited number of group check-in sequences, both group and individual check-in sequences are utilized to optimize the sequencing adapter and POI embeddings.
3.2. Enhancing Group Representations
Even with the modified LLM, it is infeasible to accurately learn the group representation from limited group check-in data. It is essential to integrate and aggregate each individual’s preferences (i.e., representation) within a group to enhance the group representation. Existing methods like averaging individual representations (Vinh Tran et al., 2019) or assigning weights based on importance or preference consistency (Yin et al., 2020) offer varying degrees of flexibility but fail to model the complex interaction patterns among group members.
In light of this, we employ the LLM to aggregate individual representations, leveraging its ability to capture intricate relationships and interactions, leading to a more nuanced and accurate group representation. Specifically, for the group $g_{n}$, we first obtain the group check-in sequence representation $\textbf{e}_{g_{n}}$ by applying the modified LLM $\Theta_{s}$ to the group check-in sequence $\mathcal{X}_{g_{n}}$ as described in Section 3.1. Then, for each member within the group $u_{k}\in\mathcal{K}_{n}$, we generate their individual representation from their check-in sequence $\mathcal{X}_{u_{k}}$:
(7) $\textbf{e}_{u_{k}}=\Theta_{s}(\mathcal{X}_{u_{k}}).$
Hence, the representations of all individual check-in sequences are denoted as $\textbf{E}_{g_{n}}=\{\textbf{e}_{u_{1}},\textbf{e}_{u_{2}},...,\textbf{e}_{u_{K}}\}$, which will be aggregated by the LLM $\Theta$, denoted as:
(8) $\textbf{e}^{\prime}_{g_{n}}=\Theta(\textbf{E}_{g_{n}}).$
Please note that $\textbf{E}_{g_{n}}$ is the input for the first transformer decoder of $\Theta$ without word embedding and positional encoding. This is because $\textbf{E}_{g_{n}}$ is a set of user representations instead of words, and the order of users is not important for preference aggregation. The output of the final transformer decoder will be utilized as the representation of the aggregated preference. Then, the enhanced group representation is obtained by:
(9) $\textbf{e}_{g_{n}}=\textbf{e}_{g_{n}}+\alpha\textbf{e}^{\prime}_{g_{n}},$
where $\alpha$ is a hyperparameter that controls the weight of the aggregated member preference.
In section 3.1, we have trained an adapter $\Theta^{{}^{\prime}}_{s}$ for generating check-in sequence representation. Obviously, this adapter cannot be applied directly to aggregate individual preferences. To distinguish between these two tasks and avoid compromising the performance of the sequencing adapter $\Theta^{{}^{\prime}}_{s}$, we introduce a new adapter $\Theta^{{}^{\prime}}_{a}$ with the same architecture as $\Theta^{{}^{\prime}}_{s}$. On this basis, Equation 8 is modified as:
(10) $\textbf{e}^{\prime}_{g_{n}}=\Theta_{a}(\textbf{E}_{g_{n}}),$
where $\Theta_{a}$ is the modified large language model by the aggregation adapter $\Theta^{{}^{\prime}}_{a}$. During the fine-tuning process, while freezing all other weights, $\Theta^{{}^{\prime}}_{a}$ is optimized by the POI prediction objective shown in Equation 6, only with the group check-in sequences and the corresponding group member check-in sequences. Here, we denote the set of all group check-in sequences and their members’ check-in sequences as $\mathcal{D}_{group}$.
Figure 5. The prompt for classifying the check-in sequence.
3.3. Self-supervised Leaning Task
As mentioned above, the LLM will be fine-tuned for check-in sequence representation generation and individual preference aggregation. However, the sparsity of user-POI interactions has been a long-standing issue in POI recommendation, which deteriorates further in group settings. To this end, we introduce a self-supervised signal to enhance representations for check-in sequences with the corresponding purposes, such as business trips and tourism, as the travel purposes provide a strong signal to predict the next POI a group/user will visit. Unfortunately, check-in data typically lacks explicit labels indicating the purpose of each check-in sequence. To overcome this limitation, we leverage the powerful capabilities of the LLM to classify each check-in sequence. The LLM analyzes the spatio-temporal and contextual information within the sequence and assigns a category that reflects its underlying purpose. To achieve this, we design the corresponding prompt which is shown in Figure 5, and we have $11$ purpose labels: {Business Trip, Work Commute, Tourism, Short Vacation, Cultural & Entertainment, Shopping & Dining, Social Events, Fitness & Wellness, Daily Commute, Medical & Health Visits, Education & Training}. The details of these labels are provided in Appendix A.1.
Beyond that, those purpose labels are only valid for short sequences, since a purposeful trip won’t last long. Therefore, we divide the whole sequence into subsequences if the time difference between any two consecutive check-in activities is larger than five days. Here, we denote the set of all short sequences as $\mathcal{D}_{ssl}$. To verify the accuracy of sequence purposes generated by the LLM, we manually label a portion ($20\%$) of short sequences as the ground truths. The LLM achieves the desirable accuracy of 92%.
To endow our model with the ability for purpose prediction, we propose a new purpose embedding matrix $\textbf{E}_{pur}\in\mathbb{R}^{d\times L}$, where $L=11$ refers to the number of purpose labels. Consequently, given a check-in sequence representation $\textbf{e}_{\mathcal{X}}$, the probability of purpose labels is calculated as:
(11) $\textbf{b}=\textbf{E}_{pur}\textbf{e}^{T}_{\mathcal{X}}.$
Then, the purpose prediction objective is defined as:
(12) $\mathcal{L}_{pur}=\mathcal{L}_{CE}(y_{p}^{{}^{\prime}},y_{p}),$
where $\mathcal{L}_{CE}$ is the cross-entropy loss, $y_{p}^{{}^{\prime}}$ is the predicted purpose by our model, and $y_{p}$ is the purpose label generated by a standard but more powerful LLM (e.g., GPT-4). This loss will be used to pre-train the sequencing adapter $\Theta^{{}^{\prime}}_{s}$ and the POI embedding matrix $\textbf{E}_{pur}$, and thus, the sparse group check-in data can leverage the relatively rich individual check-in data to obtain richer representations. The additional semantic information provided by the LLM enables the model to infer latent intents behind check-in sequences, leading to more precise and intent-aware POI recommendations.
3.4. The Optimization of LLMGPR
The optimization workflow of LLMGPR is presented in Algorithm 1. At the very beginning (line 1), we regard Llama3-8b as the base large language model $\Theta$, which can be replaced by any other large language model if sufficient computational resources are available. After that, we initialize all trainable weights including the check-in sequencing adapter $\Theta^{{}^{\prime}}_{s}$, the aggregation adapter $\Theta^{{}^{\prime}}_{a}$, the POI embedding matrix $\textbf{E}_{poi}$, and the purpose embedding matrix $\textbf{E}_{pur}$ (line 2). Then, we obtain the sequence-specific LLM $\Theta_{s}$ and the aggregation-specific LLM $\Theta_{a}$ (line 3-4). On this basis, we first pre-train the sequencing adapter $\Theta^{{}^{\prime}}_{s}$ and the POI embedding matrix $\textbf{E}_{poi}$ based on the purpose prediction loss $\mathcal{L}_{pur}(y^{\prime}_{p},y_{p})$ using all short sequences and their corresponding purpose labels (lines 5-10). The sequencing adapter $\Theta^{{}^{\prime}}_{s}$ and the POI embedding matrix $\textbf{E}_{poi}$ are further optimized by the POI prediction loss $\mathcal{L}_{seq}(y^{\prime},y)$ (lines 11-16). To address the group check-in data sparsity and obtain a better task-specific LLM for learning the representation of check-in sequences, all check-in sequences are utilized including group and all individual check-in sequences.
Subsequently, the aggregation adapter $\Theta^{{}^{\prime}}_{a}$ is trained iteratively with group check-in sequences and corresponding individual check-in sequences of each group member (lines 17-29). In detail, for each group check-in sequence, the group representation is originally obtained by applying the sequence-specific LLM $\Theta_{s}$ to the group check-in sequence (line 19). After that, individual representations for group members are obtained by applying $\Theta_{s}$ to their respective check-in sequences (lines 20-23). These representations are then fed into the aggregation-specific LLM $\Theta_{a}$ (line 24), while the aggregated representation is utilized to enhance the group representation (line 25) and generate the recommendation results (line 26). Finally, the aggregation adapter is updated using the POI prediction loss $\mathcal{L}_{seq}(y^{\prime},y)$ (line 27).
4. Experiments
In this section, we perform comprehensive experiments with two real-world datasets to evaluate the effectiveness of the proposed LLMGPR. Specifically, our investigation seeks to address the following research questions:
RQ1: How effective is LLMGPR for group POI recommendations?
RQ2: How does LLMGPR perform in learning check-in sequence representation compared to advanced POI recommender systems?
RQ3: Is each key component in LLMGPR effective?
RQ4: Does LLMGPR work well for cold-start recommendations?
RQ5: What is the influence of LLMGPR’s key hyperparameters?
4.1. Datasets and Evaluation Protocols
We adopt two real-world datasets including Foursquare (Chen et al., 2020a) and Weeplace (Liu et al., 2013) to evaluate the proposed LLMGPR, where both datasets include check-in histories in the city of New York. Since both datasets do not originally contain group information, we follow the widely adopted procedure (Chen et al., 2022) to construct group interactions by finding overlaps in check-in times and social relations. That is, we assume if a set of users who are connected on the social network visit the same venue at the same time, then they are regarded as members of a group, and the corresponding activities are group activities. The use of explicit social connections and spatiotemporal tags ensures the quality of discovered user groups. Table 5 summarizes the statistics of the two datasets.
For evaluation, we adopt the leave-one-out protocol which is widely used in previous works (Wang et al., 2018, 2019; Zheng et al., 2016). That is, for each of the check-in sequences, the last check-in POI is for testing, the second last POI is for validation, and all others are for training. In addition, the maximum sequence length is set to 200. For each ground truth check-in POI, unlike the E-commerce setting that ranks all products (Krichene and Rendle, 2020), we only pair it with 500 unvisited and nearest POIs within the same region as the candidates for ranking. The rationale is, different from e-commerce products (Krichene and Rendle, 2020; Qu et al., 2021), in the scenario of POI recommendations that are location-sensitive, users seldom travel between two POIs consecutively that are far away from each other (Long et al., 2023; Yuan et al., 2023; Li et al., 2021). On this basis, we leverage two ranking metrics, namely Hit Ratio at Rank $k$ (HR@$k$) and Normalized Discounted Cumulative Gain at Rank $k$ (NDCG@$k$) (Weimer et al., 2007) where HR@$k$ only measures the times that the ground truth is present on the top-$k$ list, while NDCG@$k$ cares whether the ground truth can be ranked as highly as possible.
For hyperparameters, the dimension size of POI embeddings is set to 4096, which is the same as that of original token embeddings in Llama3-8b. With respect to QLoRA, the quantization bit width $b$ is set to $4$ and the rank $r$ is set to $16$. Besides, we set $\alpha$ to 0.7, the learning rate to $0.0002$, the dropout to $0.2$, the batch size to $16$, and the maximum training epoch to $5$. The impacts of $r$ and $\alpha$ will be further discussed in Section 4.7.
4.2. Baselines
To evaluate the effectiveness of LLMGPR for group POI recommendations, we first compare it with existing group recommenders: MF-AVG (Vinh Tran et al., 2019), AGREE (Cao et al., 2018), GroupIM (Cao et al., 2018), CubeRec (Chen et al., 2022), HHGR (Chen et al., 2022) and MICL (Li and Zhu, 2024). For a fair comparison, we regard check-in sequence representations and POI representations generated by the well-trained LLM as initial representations of groups, users, and POIs for the above baselines. To validate the effectiveness of LLMGPR in learning check-in sequence representations, we also compare it with SOTA POI recommenders for individuals: GeoMF (Lian et al., 2014), STAN (Luo et al., 2021), DRAN (Wang et al., 2022), Diff-POI (Qin et al., 2023), META ID (Huang et al., 2024), LLMMove (Feng et al., 2024) and LLM4POI (Li et al., 2024). The details of those baselines are provided in Appendix A.4.
Table 1. Recommendation performance comparison with POI recommenders for groups.
Foursquare Weeplace
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
MF-AVG 0.2560 0.1480 0.3666 0.2143 0.2597 0.1501 0.3896 0.2320
AGREE 0.2825 0.1590 0.4003 0.2327 0.2713 0.1614 0.4226 0.2542
GroupIM 0.2917 0.1663 0.4060 0.2467 0.2852 0.1678 0.4459 0.2511
CubeRec 0.3012 0.1683 0.4170 0.2509 0.2926 0.1714 0.4567 0.2584
HHGR 0.3086 0.1795 0.4391 0.2585 0.3033 0.1825 0.4637 0.2710
MICL 0.3101 0.1797 0.4516 0.2653 0.3121 0.1834 0.4693 0.2766
LLMGPR 0.3198 0.1926 0.4875 0.2874 0.3217 0.1917 0.4978 0.2934
4.3. Group Recommendation Effectiveness (RQ1)
The performance comparison among all the group recommenders is summarized in Table 1, revealing several key insights. Firstly, AGREE consistently outperforms MF-AVG on both datasets, due to its attention mechanism that dynamically assigns weights for preference aggregation, thus better capturing the interactions and relationships among group members. However, the use of point embeddings limits its flexibility and resilience in representing groups. To address this, GroupIM and CubeRec introduce hypercube-based group representations, replacing point embeddings with range embeddings, which achieve better recommendation accuracy. Additionally, HHGR and MICL utilize graph neural networks to capture relationships between users, groups, and POIs, resulting in more precise group representations and more accurate recommendations. Despite the advances of these methods, the proposed LLMGPR model demonstrates superior accuracy over all baselines. Specifically, compared to the best-performing baseline, MICL, LLMGPR achieves average improvements of 5.05% on Foursquare and 4.86% on Weeplace. Another notable observation is that all the group recommenders achieve higher recommendation accuracy on Weeplace than on Foursquare, due to its sparser group-POI interactions and a larger set of POIs. Notably, the proposed LLMGPR can still leverage augmented supervision signals to learn quality group representations, achieving the best group recommendation performance even on this challenging dataset.
Table 2. Recommendation performance comparison with POI recommenders for individuals.
Foursquare Weeplace
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
GeoMF 0.1051 0.0604 0.1326 0.0892 0.0849 0.0610 0.0958 0.0664
STAN 0.3113 0.1801 0.4654 0.2896 0.3004 0.1761 0.4361 0.2621
DRAN 0.3159 0.1826 0.4816 0.2970 0.3107 0.1807 0.4349 0.2677
Diff-POI 0.3211 0.1935 0.4870 0.2918 0.3106 0.1848 0.4751 0.2827
META ID 0.3058 0.1799 0.4711 0.2915 0.3075 0.1813 0.4338 0.2702
LLMMove 0.3146 0.1892 0.4737 0.2813 0.3062 0.1829 0.4763 0.2763
LLM4POI 0.3241 0.1996 0.4976 0.2970 0.3133 0.1882 0.4898 0.2936
LLMGPR 0.3386 0.2129 0.5101 0.3037 0.3284 0.2027 0.5072 0.3012
4.4. Effectiveness of Learning Check-in Sequence Representations (RQ2)
To further study the effectiveness of the proposed LLMGPR in learning check-in sequence representation, we also compare LLMGPR with SOTA POI recommenders for individuals, where the results are summarized in Table 2. Our proposed LLMGPR consistently and significantly beats the baselines on both datasets. As the basic POI recommendation approach without considering the sequential effect of user check-in activities, GeoMF undoubtedly has the worst performance on both datasets. STAN outperforms GeoMF by a significant margin since it can leverage spatiotemporal correlations in both consecutive and non-consecutive check-in activities with the attention mechanism. Advancing further, DRAN melds a Graph Neural Network (GNN) with an attention mechanism, leading to more refined POI embeddings and thus outperforming STAN in terms of accuracy. Most notably, Diff-POI, employing the robust generality of the diffusion model, sets a new benchmark for state-of-the-art accuracy in this domain.
Recent works focus on integrating the LLM for recommendations, due to its ability to leverage the semantic information of users and items. META ID uses out-of-vocabulary tokens to characterize users and items for recommendations. As it is not designed for POI recommendations, it cannot integrate and exploit the spatial and temporal information of check-ins, even performing worse than Diff-POI. In addition, LLMMove applies the zero-shot generation ability of the LLM for the next POI recommendations. Without fine-tuning for the specific recommendation task, its performance is limited. Instead, LLM4POI fine-tunes the LLM on the two datasets for the POI recommendation task, and thus, it achieves better performance. Notably, our proposed LLMGPR significantly outperforms the best baseline LLM4POI, showing its superiority in learning check-in sequence representations. This is attributed to the well-trained sequencing adapter and the introduction of a novel self-supervised learning task.
Table 3. Ablation study for LLMGPR.
Foursquare Weeplace
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
LLMGPR 0.3198 0.1926 0.4875 0.2874 0.3217 0.1917 0.4978 0.2934
LLMGPR-FT 0.2781 0.1652 0.4159 0.2355 0.2947 0.1704 0.4312 0.2439
LLMGPR-M 0.2834 0.1688 0.4367 0.2495 0.2912 0.1704 0.4472 0.2592
LLMGPR-ER 0.2921 0.1701 0.4332 0.2524 0.2945 0.1719 0.4513 0.2666
LLMGPR-SSL 0.3119 0.1816 0.4623 0.2781 0.3091 0.1825 0.4732 0.2847
4.5. Ablation Study (RQ3)
In this section, we perform the ablation analysis to evaluate the performance gain from the core components of LLMGPR in terms of the main task: group POI recommendations. Table 3 summarizes the recommendation outcomes for different degraded versions of LLMGPR. Then, we introduce all variants and discuss the effectiveness of corresponding model components.
Using the LLM directly without fine-tuning (LLMGPR-FT). In LLMGPR, we modify and fine-tune the LLM for group POI recommendations. To evaluate the usefulness of those modifications, we apply the original LLM directly for group POI recommendations, where both group-level and member-level check-in sequences are utilized. This variant model has experienced noticeable performance drops, especially on the Foursquare dataset with the higher sparsity of group-item interactions. Hence, including modification and fine-tuning of the LLM is important for LLMGPR to learn representative group representations in the scenario of data sparsity.
Using the fine-tuned but not modified LLM (LLMGPR-M). This variant differs from LLMGPR-FT as it is fine-tuned using the QLoRA strategy. It can also be seen as directly applying LLM4POI (Li et al., 2024) for group POI recommendations. As anticipated, the recommendation accuracy drops significantly, underscoring the effectiveness of the additional POI embedding matrix and the innovative approach to constructing group representations.
Removing the enhancement of group representation with aggregated member preferences (LLMGPR-ER). A crucial component of LLMGPR is enhancing group representation by aggregated member preference as described in Section 3.2. To test its effectiveness, we remove it and only utilize the group check sequence for final recommendations. As a result, there is a significant decrease in recommendation accuracy for both datasets. This reflects that, even with the well-trained LLM, the sparse group-item interactions significantly harm the accuracy, and our proposed aggregation strategy can effectively address this issue.
Removing self-supervised learning (LLMGPR-SSL). This variant disables the self-supervised loss when pretraining the sequencing adapter $\Theta^{{}^{\prime}}_{s}$ and the POI embedding matrix $\textbf{E}_{poi}$, as shown in Section 3.3. As LLMGPR can no longer leverage the purposes of short sequences as a supervision signal to counteract the group-level data sparsity, it suffers from inferior recommendation accuracy on both datasets. Therefore, the novel self-supervised learning scheme showcases its strong contribution to enriching group-level information.
4.6. Effect of Cold-start Recommendations (RQ4)
To demonstrate LLMGPR’s ability in cold-start recommendations, we randomly selected $200$ group-level check-in sequences from each dataset. These sequences were excluded from the training process and each had fewer than 10 check-ins. We then recorded the average recommendation accuracy (HR@10) for these sequences across all group POI recommenders, including LLMGPR, as shown in Table 4. LLMGPR’s cold-start accuracy was notably higher than all baselines, highlighting its effectiveness in providing recommendations for new groups with limited historical check-in data.
Table 4. Cold-start recommendation performance (HR@10).
Foursquare Weeplace
MF-AVG 0.3148 0.3203
AGREE 0.3605 0.3537
GroupIM 0.3698 0.3849
CubeRec 0.3811 0.3925
HHGR 0.4178 0.4274
MICL 0.4157 0.4192
LLMGPR 0.4689 0.4663
4.7. Hyperparameter Sensitivity (RQ5)
In this section, we illustrate the effect of two hyperparameters on the group-level recommendation accuracy of LLMGPR including the rank $r$ which controls adapter sizes, and $\alpha$ that controls the injection level of the aggregated members’ preference to the group representations. The results are shown in Figure 6.
Impact of $r$. Recommendation accuracy is recorded for $\lambda\in\{4,8,16,32,64\}$. Usually, the recommendation accuracy benefits from higher $r$. However, the accuracy stabilizes once $r$ exceeds $16$, indicating that LLMGPR is capable of delivering high-performance recommendations without requiring additional resources to train larger adapters.
Impact of $\alpha$. We experiment on $\alpha\in\{0,0.1,0.3,0.5,0.7,0.9,1\}$. The lowest accuracy is obtained if the aggregated members’ preference is not applied ($\alpha=0$), showing the significance of the strategy for enhancing group representations. At the start, the recommendation accuracy increases with the increase of $\alpha$. However, the accuracy will decline if knowledge of members’ individual preferences has an excessive proportion ($\alpha>0.7$).
Figure 6. Hyperparameter sensitivity.
5. Related work
This section reviews recent literature on related areas including POI recommendations and group POI Recommendations.
5.1. POI Recommendations
To help people discover attractive places by analyzing user-POI interactions, early works on discovering attractive places through user-POI interactions primarily focused on matrix factorization (Lian et al., 2014) and Markov chains (Cheng et al., 2013; Zheng et al., 2016). The advent of recurrent neural network (RNN) models marked a significant advancement, as they demonstrated the ability to capture spatiotemporal dynamics in POI sequences (Chen et al., 2020b; Jie et al., 2018; Chen et al., 2020c; Yin et al., 2015; Li et al., 2021). Additionally, models utilizing attentive neural networks (Luo et al., 2021; Yang et al., 2022; Chen et al., 2019; Yin et al., 2015) adopted self-attention mechanisms to analyze the spatiotemporal context of sequential check-ins. Graph neural network (GNN)-based models (Li et al., 2021; Wang et al., 2022; Gao et al., 2023a, b) further advanced the field by integrating graph-augmented POI sequences, leveraging collaborative signals from semantically similar POIs to reveal sequential trends and outperform RNN-based approaches in accuracy. Diff-POI (Qin et al., 2023) sets a new standard by using diffusion models to achieve cutting-edge accuracy. Recent research focuses on integrating Large Language Models (LLMs) for POI recommendations. LLMMove (Feng et al., 2024) employs zero-shot generation for next POI recommendations, while LLM4POI (Li et al., 2024) fine-tunes LLMs on related datasets to exploit commonsense knowledge. Compared to these approaches, LLMGPR demonstrates significant advantages in learning sequence representations and addressing group-level data sparsity by optimizing adapters and the POI embedding matrix.
5.2. Group POI Recommendations
Group POI recommenders typically involve two main phases: learning group representations from group-level check-in sequences and enhancing these representations using aggregated member representations derived from individual check-in sequences. Existing methods for preference aggregation range from simple averaging (Vinh Tran et al., 2019), which treats all members equally, to weighted averaging (Sojahrood and Taleai, 2021), which assigns weights based on external factors but struggles with determining appropriate values. More sophisticated approaches, such as attentive neural networks (Cao et al., 2018; Guo et al., 2020), hypercube-based embeddings (Sankar et al., 2020; Chen et al., 2022), and graph convolution networks (Zhang et al., 2021; Li and Zhu, 2024), aim to capture multifaceted preferences and complex relationships among users, groups, and POIs. However, these models often face challenges like overfitting and lack of scalability. In contrast, LLMGPR, a modified and well-trained LLM, can capture complex patterns and adapt dynamically to individual preferences, offering improved performance and adaptability for group POI recommendations.
5.3. LLM-based Recommendations
Motivated by the notable achievements of LLMS, the integration of LLMs into recommendation systems has grown significantly. Early works (Feng et al., 2024; Hou et al., 2024) primarily focused on designing prompts to utilize LLMS directly with zero-shot generalization. However, these approaches often struggled with the lack of specificity and domain adaptation, which led to the development of prompt tuning techniques (Zhai et al., 2023; Wu et al., 2024). Prompt tuning involved adjusting the prompts to better suit the recommendation task, offering improved performance but still facing limitations in handling complex recommendation scenarios. This led to the exploration of fine-tuning LLMs (Lin et al., 2024; Bao et al., 2023), where models were adapted to the recommendation tasks through task-specific training, significantly enhancing accuracy but lacking unique indicates of item IDs. Hence, the construction of IDs was introduced (Huang et al., 2024; Yu et al., 2024), creating embeddings or representations that could capture nuanced user preferences and item characteristics, thus offering a more efficient and scalable solution to recommendation problems. LLMGPR introduces and optimizes additional weights, including POI embeddings, sequencing adapters, and aggregation adapters, to modify the LLM for group-level recommendations. These modifications aim to learn sequence representations and aggregate members’ preferences more effectively.
6. Conclusion
In this paper, we introduced LLMGPR, a novel large language model-based framework designed to address the challenges of group POI recommendation in Location-Based Social Networks. Our approach effectively learns sequence representations and dynamically aggregates group members’ preferences, overcoming the limitations of existing methods that struggle with data sparsity and inadequate preference aggregation. By leveraging the capabilities of large language models and introducing additional POI tokens, and adapters for sequencing and aggregation tasks, LLMGPR generates dense, holistic representations that improve recommendation accuracy. Additionally, a self-supervised learning signal enhances performance and generalizability across diverse groups. Our experiments demonstrate significant improvements over existing methods, highlighting the potential of LLMGPR to revolutionize group POI recommendations and better satisfy the diverse preferences of group members.
Appendix A Appendix
A.1. Short sequence purple labels
All short sequence labels for self-supervised learning in Section 3.3 are:
  • Business Trip: This label captures check-ins related to work travel, such as meetings, conferences, or site visits. Business trips are often brief but involve specific destinations like hotels, airports, and office buildings, making them important for identifying professional user behaviors.
  • Work Commute: This label covers routine, daily commutes to and from work. It captures repetitive patterns in user check-ins tied to their professional lives, providing useful insights into regular, predictable movements between home and work locations.
  • Tourism: This category includes check-ins at tourist attractions, such as landmarks, museums, or vacation destinations. Tourism-related trips are typically short in duration but involve diverse POIs, making it a crucial label for understanding users’ travel behavior.
  • Short Vacation: This label captures brief leisure trips that are not necessarily tied to family or friends but are characterized by relaxation and exploration over a few days. Such vacations may include weekend getaways, short stays at resorts, or visits to nearby tourist destinations. Unlike long trips, short vacations typically involve fewer check-ins but reflect an intention to explore new locations within a limited time, providing valuable insight into recreational user behavior.
  • Cultural & Entertainment: This label covers activities like visiting theaters, concerts, galleries, or other entertainment venues. Cultural outings often represent unique or occasional check-ins, distinct from routine patterns, and are driven by leisure and enjoyment.
  • Shopping & Dining: This category encompasses visits to shopping centers, retail stores, and restaurants. These activities are typically short-term, common in urban exploration, and represent essential behaviors in users’ leisure and daily lives.
  • Social Events: This label covers participation in larger gatherings like weddings, parties, or festivals. Social events drive check-ins at specific venues and are typically tied to one-off, significant occasions.
  • Fitness & Wellness: This category includes gym visits, outdoor runs, or wellness centers. These check-ins often happen regularly within short time spans, reflecting health-focused user habits.
  • Daily Commute: This label captures non-work-related daily travel, such as trips to grocery stores, schools, or other essential locations. These are habitual, short-duration check-ins critical for understanding routine user behavior outside of work.
  • Medical & Health Visits: This label covers check-ins at hospitals, clinics, or pharmacies. These visits are generally short-term and occur due to specific health needs, adding an important dimension to occasional user behavior.
  • Education & Training: This category includes check-ins related to learning environments, such as schools, universities, or training centers. These trips often involve short-term attendance but are crucial for understanding education-driven travel.
These labels are designed to reflect the most common and distinct purposes for short-term trips or sequences of check-ins. Each category is tied to a specific motivation or context that drives user mobility, ranging from daily routines to more occasional events. Together, they comprehensively capture the various reasons behind short-duration check-ins, ensuring that each sequence can be classified meaningfully and relevantly for POI prediction.
A.2. Algorithm 1: The optimization of LLMGPR
Algorithm 1 The optimization of LLMGPR.
1:/*Pretrain weights for sequence representation generation */
2:Obtain the large language model $\Theta$;
3:Initialize $\Theta^{{}^{\prime}}_{s}$, $\Theta^{{}^{\prime}}_{a}$, $\textbf{E}_{poi}$, $\textbf{E}_{pur}$;
4:Obtain $\Theta_{s}$ by combining $\Theta$ and $\Theta^{{}^{\prime}}_{s}$;
5:Obtain $\Theta_{a}$ by combining $\Theta$ and $\Theta^{{}^{\prime}}_{a}$;
6:repeat
7:   for $(\mathcal{X},y_{p})\in\mathcal{D}_{ssl}$ do
8:      $y^{\prime}_{p}\leftarrow\Theta_{s}(\mathcal{X},\textbf{E}_{poi},\textbf{E}_{pur})$;
9:      Take a gradient step w.r.t $\mathcal{L}_{pur}(y^{\prime}_{p},y_{p})$ to update $\Theta^{{}^{\prime}}_{s}$, $\textbf{E}_{poi}$, $\textbf{E}_{pur}$;
10:   end for
11:until convergence
12:/*Train weights for sequence representation generation */
13:repeat
14:   for $(\mathcal{X},y)\in\mathcal{D}$ do
15:      $y^{\prime}\leftarrow\Theta_{s}(\mathcal{X},\textbf{E}_{poi})$;
16:      Take a gradient step w.r.t $\mathcal{L}_{seq}(y^{\prime},y)$ to update $\Theta^{{}^{\prime}}_{s}$, $\textbf{E}_{poi}$;
17:   end for
18:until convergence
19:/*Train the aggregation adapter*/
20:repeat
21:   for $(\mathcal{X}_{g_{n}},\mathcal{K}_{n},y)\in\mathcal{D}_{group}$ do
22:      $\textbf{e}_{g_{n}}=\Theta_{s}(\mathcal{X}_{g_{n}},\textbf{E}_{poi})$
23:      for $\mathcal{X}_{u_{k}}\in\mathcal{X}(\mathcal{K}_{n})$ do
24:         $\textbf{e}_{u_{k}}=\Theta_{s}(\mathcal{X}_{u_{k}},\textbf{E}_{poi})$;
25:      end for
26:      $\textbf{E}_{g_{n}}=\{e_{u_{1}},e_{u_{2}},...,e_{u_{K}}\}$;
27:      $\textbf{e}^{\prime}_{g_{n}}=\Theta_{a}(\textbf{E}_{g_{n}})$;
28:      $\textbf{e}_{g_{n}}=\textbf{e}_{g_{n}}+\alpha\textbf{e}^{\prime}_{g_{n}}$;
29:      $y^{\prime}\leftarrow\textbf{E}_{poi}(\textbf{e}_{g_{n}})$;
30:      Take a gradient step w.r.t $\mathcal{L}_{seq}(y^{\prime},y)$ to update $\Theta^{{}^{\prime}}_{a}$;
31:   end for
32:until convergence
A.3. Dataset statistics.
Table 5. Dataset statistics.
Foursquare Weeplace
#users 6,078 3,795
#groups 1,557 867
#POIs 63,445 35,911
#categories 436 625
#user check-ins 923,856 519,915
#group check-ins 10,899 10,404
#check-ins per user 152 137
#check-ins per group 7 12
#users per group 3.67 4.58
A.4. Baselines
Group POI Recommenders:
  • MF-AVG (Vinh Tran et al., 2019): It averages group-wise user representations learned via user-item matrix factorization to form group embeddings.
  • AGREE (Cao et al., 2018): It incorporates an attention mechanism to deduce group preferences from individual member preferences and aggregates them into a unified group preference vector.
  • GroupIM (Sankar et al., 2020): It Utilizes expressive hypercubes in vector space with a new distance metric and an intersection-based self-supervision paradigm to facilitate group-item pairwise ranking.
  • CubeRec (Chen et al., 2022): It utilizes expressive hypercubes in vector space with a new distance metric and an intersection-based self-supervision paradigm to facilitate group-item pairwise ranking.
  • HHGR (Zhang et al., 2021): It employs a hypergraph neural network to encapsulate complex high-order relationships among users, groups, and items, with a double-scale self-supervised learning objective.
  • MICL (Li and Zhu, 2024): It employs the group-view adaptive graph transformer, member-view hypergraph aggregation network, and item-view tripartite graph augmentation to deeply explore potential interactive compromises.
Normal POI Recommenders:
  • MF (Lian et al., 2014): It utilizes user-item matrix factorization for POI recommendations.
  • STAN (Luo et al., 2021): It learns explicit spatiotemporal correlations of check-in trajectories via a bi-attention approach.
  • DRAN (Wang et al., 2022): It leverages a disentangled representation-enhanced attention network for next POI recommendation using GNN-based methods.
  • Diff-POI (Qin et al., 2023): It uses a diffusion-based model to sample from the posterior distribution that reflects user geographical preferences.
  • META ID (Huang et al., 2024): It introduces ID representations in LLM-based recommender systems, and uses out-of-vocabulary tokens to characterize users/items, enhancing the memorization and diversity of their representations.
  • LLMMove (Feng et al., 2024): It applies zero-shot generalization of the LLM for next POI recommendations.
  • LLM4POI (Li et al., 2024): It fine-tunes the LLM on standard-sized datasets to exploit commonsense knowledge for the next POI recommendations.